### v1

# Breaking Annotation Barriers: Generalized Video Quality Assessment via Ranking-based Self-Supervision

Abstract
Video quality assessment (VQA) is essential for quantifying perceptual quality in various video processing workflows, spanning from camera capture systems to over-the-top streaming platforms. While recent supervised VQA models have made substantial progress, the reliance on manually annotated datasets—a process that is labor-intensive, costly, and difficult to scale up—has hindered further optimization of their generalization to unseen video content and distortions. To bridge this gap, we introduce a self-supervised learning framework for VQA to learn quality assessment capabilities from large-scale, unlabeled web videos. Our approach leverages a learning-to-rank paradigm to train a large multimodal model (LMM) on video pairs automatically labeled via two manners, including quality pseudo-labeling by existing VQA models and relative quality ranking based on synthetic distortion simulations. Furthermore, we introduce a novel iterative self-improvement training strategy, where the trained model acts an improved annotator to iteratively refine the annotation quality of training data. By training on a dataset 10×10$\times$10 × larger than the existing VQA benchmarks, our model: (1) achieves zero-shot performance on in-domain VQA benchmarks that matches or surpasses supervised models; (2) demonstrates superior out-of-distribution (OOD) generalization across diverse video content and distortions; and (3) sets a new state-of-the-art when fine-tuned on human-labeled datasets. Extensive experimental results validate the effectiveness of our self-supervised approach in training generalized VQA models. The datasets and code will be publicly released to facilitate future research.
Figure 1: Overview of our work. (a) Existing state-of-the-art VQA models exhibit poor out-of-distribution performance. (b) We construct a large-scale VQA dataset consisting of $700$k video pairs, sampled from multiple social media platforms, covering more than 20 content categories. (c) We explore two strategies for automatically annotating the relative quality of video pairs. (d) Our proposed model undergoes iterative training to enhance generalization performance.
footnotetext: *Equal contribution. Project leader.
1 Introduction
Video quality assessment (VQA) [32] This work focuses on no-reference (NR) or blind VQA, which assesses video quality without relying on additional reference information. plays an important role in modern video processing systems, delivering objective quality measurements used to optimize end-user Quality of Experience (QoE). With the advances in deep neural networks (DNNs) [12, 7, 26] and the increasing availability of human-annotated VQA datasets [13, 47, 56, 65], current VQA models [49, 57, 60, 61] have achieved significant progress through supervised learning. Nevertheless, supervised learning inherently faces a limitation: the generalization of the VQA models heavily depends on the diversity of the training data. For example, even top-tier VQA models [49, 57, 60, 61] exhibit significant performance drops in out-of-distribution evaluations, as illustrated in Fig. 1(a).
Therefore, the development of specialized VQA datasets remains critical to address the growing diversity of emerging media formats and their associated distortions [70, 69, 46, 63, 23, 30, 66]. However, constructing such datasets is highly resource-intensive. A standardized subjective experiment comprises two key phases: test sample curation and subjective quality annotation. The test sample curation phase necessitates rigorous selection of representative video samples, as inadequate sampling strategies risk producing oversimplified datasets (i.e.,easy dataset” problem [49, 3]) and may induce model overfitting. Meanwhile, subjective annotation—though vital—is laborious and costly. International Telecommunication Union (ITU) standards (e.g., ITU-T P.910 [15]) outline specific recommendations for experimental setups, including display conditions, stimulus duration, subject count, and rating methodologies. These constraints, though necessary for statistically meaningful results, impede large-scale dataset expansion due to prohibitive annotation costs.
Self-supervised or unsupervised learning [10], which eliminates the need for human-labeled annotations, is a potential solution in VQA to mitigate the high costs of subjective experiments. However, current self-supervised VQA methods [5, 4, 6, 31, 33] still lag behind their supervised counterparts in performance. Typical implementations of current self-supervised VQA methods [4, 31] employ contrastive learning frameworks with proxy tasks such as distortion type/severity classification on synthetically generated data. However, it suffers from two shortcomings: (1) they fail to capture visual content and aesthetic characteristics relevant to perceptual quality assessment, and (2) they inadequately model real-world authentic distortion patterns that follow complex nonlinear degradation processes.
In this paper, we propose a novel self-supervised learning framework for VQA to address the aforementioned challenges. Specifically, we reformulate quality regression as a ranking problem, enabling the model to learn quality assessment capabilities through pairwise comparisons. This approach is motivated by the observation that pairwise ranking is more intuitive and reliable than absolute quality ratings for human evaluators. We then explore two strategies to automatically label the relative quality of video pairs. The first leverages existing VQA model as the “judge” to determine the relative quality, where we ensemble the results from multiple judges to mitigate inherent evaluation biases in individual models. This process is iterative: once a new VQA model is trained, it can serve as an enhanced judge, achieving self-improvement through repeated refinement. The second approach relies on synthetic distortion simulations, where we introduce various types of distortions at different severity levels and utilize these severity levels to establish relative quality rankings.
We adopt a large multimodal model (LMM) integrated with a fixed motion encoder and a learnable motion projector as our VQA model. Given two videos and a text prompt as input, the model outputs a textual judgment indicating whether the first video has higher or lower quality than the second. We employ a standard supervised fine-tuning (SFT) strategy to train the proposed VQA model using automatically annotated video pairs. Training is conducted in three iterative stages with progressively increasing sample sizes of $500$k, $600$k, and $700$k video pairs, respectively. We evaluate the proposed model on ten diverse VQA benchmarks. Experimental results show that the model achieves zero-shot performance comparable to, or even surpassing, existing supervised VQA models, while demonstrating superior out-of-distribution generalization. Moreover, after fine-tuning on human-annotated datasets, the model significantly outperforms state-of-the-art supervised approaches.
Our key contributions are summarized as follows:
  • We construct a large-scale VQA dataset comprising 700k video pairs with both authentic and synthetic distortions, each annotated with automatically generated quality labels.
  • We propose a self-supervised VQA framework that enables the model to learn quality assessment capabilities through pairwise comparisons, enhanced by an iterative training strategy for continuous self-improvement.
  • Our model achieves strong zero-shot performance across multiple VQA benchmarks, highlighting its effectiveness and generalization in video quality assessment.
2 Related Work
2.1 VQA Models
Supervised VQA. Early supervised VQA models, such as V-BLIINDS [43], TLVQM [19], and VIDEVAL [52], are typically knowledge-driven. These methods extract quality-related features based on natural scene statistics (NSS) [34], blurriness [19], motion vectors [18], optical flow [2], and other handcrafted cues, followed by training a machine learning-based regressor, such as a support vector regressor (SVR) or random forest regressor, to derive quality scores.
With the popularity of DNNs, some VQA methods, such as VSFA [22], Li22 [20], PatchVQA [65], etc., employ pre-trained DNNs as feature extractors to derive quality representations from video frames, followed by training quality regressors to map the extracted features into quality scores. Commonly used feature extractors include image classification models [12], image quality assessment (IQA) models [68, 64], and action recognition models [9], while commonly used quality regressors often consist of GRUs [22], Transformers [58], and InceptionTime [65], etc.
Many recent works explore the use of 2D or 3D convolutional neural networks (CNNs) [12], Vision Transformers (ViTs) [7], and LMMs [61] as feature extractors, fine-tuning them in an end-to-end manner to achieve state-of-the-art performance. For instance, SimpleVQA [48] and MinimalisticVQA [49] perform end-to-end training of the spatial feature extractor (e.g., ResNet [12], Swin [26]) while adopting a temporal extractor (i.e., SlowFast [9]) with fixed pre-trained weights. FAST-VQA [57] trains a 3D DNN (i.e., VideoSwin [28]) in an end-to-end fashion, and DOVER [60] further extends FAST-VQA with an aesthetic evaluation branch based on a 2D CNN (i.e., ConvNet [27]). Q-Align [61] fine-tunes the visual encoder of an LMM with a predefined text-based quality inquiry prompt.
Unsupervised and self-supervised VQA. A class of popular unsupervised or self-supervised VQA approaches [4, 31, 33, 62] aims to learn quality-aware feature representations from scratch, followed by fine-tuning a linear projector with human-annotated labels to serve as a quality score regressor. For example, CSPT [4] and CONVIQT [31] adopt contrastive learning frameworks with proxy tasks such as next-frame feature discrimination and distortion type/severity classification to learn the quality-aware representation. QPT V2 [62], on the other hand, employs an encoder-decoder architecture to reconstruct pristine videos from distorted ones, thereby learning distortion-aware representations.
Another line of research focuses on developing opinion-unaware VQA methods, which aligns with the objective of this work. Some knowledge-driven VQA methods [35, 67, 24, 36, 17] estimate video quality by directly measuring the distribution distance of specific features between distorted and pristine videos. For instance, NIQE [35] assesses the spatial quality of a test image by computing the distance between the Multivariate Gaussian model fitted to local features of the test image and pristine natural images. TPQI [24] captures temporal distortions by analyzing the straightness and compactness of video trajectories within perceptual domains of the human visual system. Recent studies [60, 1] leverage visual-language models to achieve zero-shot I/VQA. For example, BUONA-VISTA [59] employs CLIP [41] to estimate the relatively probabilities of text promt “high quality” compared to “low quality”, and combines these with NIQE and TPQI scores to assess video quality.
2.2 VQA Datasets
The progress of data-driven VQA models heavily relies on VQA datasets annotated by human subjects. Early VQA datasets, such as LIVE-VQA [45] and CSIQ-VQA [54], primarily focus on compression and transmission distortions in professionally generated content (PGC). However, these datasets contain a limited number of source videos and only a few hundred distorted samples. While they have contributed to the development of knowledge-driven VQA methods, their scale and diversity make them inadequate for training data-driven VQA models. With the rise of social media platforms, the focus has shifted to user generated content (UGC) videos characterized by natural and in-the-wild distortions. Current mainstream UGC VQA datasets, such as KoNVid-1k [13], YouTube-UGC [56], and LSVQ [65], contain thousands to tens of thousands of videos, providing large-scale benchmarks that have significantly advanced research in data-driven VQA. Meanwhile, with the emergence of new media formats, specialized VQA datasets have been developed to address domain-specific challenges, such as gaming videos [66], 360-degree videos [63], 4K videos  [23], high-frame-rate videos [30], AIGC videos [69], etc., each designed to facilitate the development of expert VQA models tailored to specific video types.
3 Pairwise-Labeled Video Dataset
This section describes the construction of the large-scale pairwise-labeled video dataset (PLVD) in detail.
Figure 2: Distribution of nine metrics across mainstream UGC datasets (LSVQ, KoNVid-1k, YouTube-UGC, LIVE-VQC), as well as our dataset before and after sampling.
3.1 Video Collection
Source Video Collection. We create a large-scale dataset comprising $3$ million videos collected from popular social media platforms, including YouTube, TikTok, Youku, and Bilibili. This dataset encompasses a wide range of content categories and distortion scenarios, such as vlogs, gaming videos, animations, live streams, etc., ensuring a representative and diverse collection with varying quality conditions. We provide the detail analysis of source videos in the supplement file.
Candidate Video Sampling. We select a subset of videos from the collected pool using a mixed-integer programming approach [53] to match target distributions defined by nine low-level metrics that quantify visual characteristics of datasets, including blockiness [42], blur [38], contrast [40], noise, flickering [39], colorfulness [11], luminance, temporal information (TI) [16], and spatial information (SI) [16]. We provide the calculation of these metrics in the supplement file. Our target distribution mirrors the aggregated distributions of mainstreaming UGC datasets—LSVQ [65], KonViD-1k [13], YouTube-UGC [56], and LIVE-VQC [47]—to ensure compatibility with existing benchmarks. Finally, we sample $438$k videos to enhance diversity in content and scenes while maintaining close alignment with mainstream UGC datasets. As illustrated in Fig. 2, the distributions of the sampled subset exhibit strong consistency with that of public VQA datasets across all nine metrics.
3.2 Pairwise Quality Annotation
We explore two strategies to automatically annotate the relative quality of video pairs: (1) quality pseudo-labeling using existing VQA models, and (2) relative quality ranking based on synthetic distortion simulations. We define two types of ranking labels: hard ranking and soft ranking. Specifically, hard ranking consists of two categories, “better” and “worse”, while soft ranking provides finer-grained comparisons with five levels: “superior”, “better”, “similar”, “worse”, and “inferior”, to reflect varying degrees of relative quality.
Table 1: Statistics of raw videos and video pairs in the PLVD dataset. The values in each cell indicate the number of videos or video pairs in PLVD-Part1/-Part2/-Part3
Videos Video Pairs
PLVD-VQA 200k/100k/50k 250k/85k/85k
PLVD-DS Spatial 50k/2k/2k 160k/5k/5k
Temporal 20k/1k/1k 40k/5k/5k
Compression 10k/1k/1k 50k/5k/5k
Total 280k/384k/438k 500k/600k/700k
3.2.1 Pseudo-labeling using VQA Models
Inspired by recent LLM-as-a-Judge methods [55] that leverage large language models (LLM) to model user preference for LLM alignment, we propose utilizing established VQA models as judges to evaluate the relative quality of video pairs. We refer to this subset of data as PLVD-VQA. Specifically, we choose five state-of-the-art VQA models: Minimalistic-VQA (VII) [49], Minimalistic-VQA (IX) [49], FAST-VQA [57], DOVER [60], and Q-Align [61], all trained on LSVQ [65], as our initial judges. For a video pair $(x^{A},x^{B})$, each judge model $\mathcal{J}_{i}$ generates quality scores $j^{A}_{i}$ and $j^{B}_{i}$. For hard ranking, the judgment is determined by comparing the mean quality scores $\overline{j}^{A}$ and $\overline{j}^{B}$. If $\overline{j}^{A}>\overline{j}^{B}$, we annotate $x^{A}$ as higher quality than $x^{B}$ (label: “better”); otherwise, $x^{A}$ is labeled as worse quality (label: “worse”).
For soft ranking, we further compute the score variance for the video pair $\sigma^{2}_{A}$ and $\sigma^{2}_{B}$. Assuming the quality difference $\Delta=\overline{j}^{A}-\overline{j}^{B}$ follows a Gaussian distribution $\mathcal{N}(\Delta;0,\sigma^{2}_{\Delta})$, where $\sigma_{\Delta}=\sqrt{\sigma^{2}_{A}+\sigma^{2}_{B}}$, we assign labels based on statistical significance thresholds adapted from [71]: pairs are labeled as “superior” if $\Delta>2\sigma_{\Delta}$, “better” if $\sigma_{\Delta}<\Delta\leq 2\sigma_{\Delta}$, “similar” if $-\sigma_{\Delta}<\Delta\leq\sigma_{\Delta}$, “worse” if $-2\sigma_{\Delta}<\Delta\leq-\sigma_{\Delta}$, and “inferior” if $\Delta\leq-2\sigma_{\Delta}$.
3.2.2 Quality Ranking via Distortion Simulations
Figure 3: Overall framework of the proposed LMM-PVQA model. Compared to the standard LMM, our visual feature extractor adopts a dual-branch design with an additional motion encoder and projector to capture temporal-related distortions. The model takes a video pair and a text prompt as inputs and generates a text output specifying which video has better quality.
We introduce synthetic distortions to simulate typical degradation that may occur in real-world scenarios, which are categorized into three types: spatial distortions, temporal distortions, and streaming distortions. Specifically, spatial distortions include resolution resizing, Gaussian blur, Gaussian noise, darkening, and brightening, which simulate capture-related artifacts. Temporal distortions consist of jitter and stuttering, mimicking playback issues commonly observed in practical settings. Streaming distortions involve H.264 and H.265 compression, reflecting compression artifacts introduced by streaming media platforms. We denote this subset of data as PLVD-DS. The detailed generation of synthetic distortions are provided in the supplement file.
We leverage distortion severity levels (e.g., constant rate factor for compression) as pseudo-labels to infer relative quality. Given a primary video $x^{0}$ and a synthetic distortion simulator $\mathcal{S}$, we degrade $x^{0}$ across $N_{\mathcal{S}}$ severity levels to generate distorted videos $\{x_{\mathcal{S}}^{i}\}^{N_{\mathcal{S}}}_{i=1}$. Pairs $(x_{\mathcal{S}}^{i},x_{\mathcal{S}}^{j})$ are randomly sampled. For hard ranking, pairs are directly annotated as “better” if $i<j$ or “worse” otherwise. For soft ranking, pairs with a severity difference $|i-j|>1$ are labeled as “superior” or “inferior” depending on the relative order of $i$ and $j$, while pairs with $|i-j|=1$ receive “better” or “worse”. The “similar” label is intentionally excluded, as $i-j=0$ implies identical videos.
3.2.3 Label Refinement
The label quality of the PLVD-VQA dataset inherently depends on the performance of judges. To address this dependency, we introduce an iterative label refinement framework. Specifically, once a new VQA model is trained, we treat it as an improved judge $\mathcal{J}^{\prime}$ and reapply the annotation pipeline of PLVD-VQA to iteratively refine video pair labels.
In summary, the PLVD dataset comprises $700$k annotated video pairs, partitioned into three subsets of $500$k, $100$k, and $100$k pairs, denoted as PLVD-Part1, PLVD-Part2, PLVD-Part3. The latter two subsets are dedicated to iterative label refinement via the trained model. A detailed breakdown of PLVD, including pair types and the corresponding number of videos, is provided in Table 1.
4 Proposed Method
We aim to train a VQA model on an unlabeled video dataset to compute the perceptual quality score of a video. To achieve this goal, we reformulate the VQA regression task as a classification problem that distinguishes the relative quality between pairs of videos.
4.1 Model Structure
We introduce an LMM-based VQA framework for pairwise quality ranking (LMM-PVQA). As illustrated in Fig. 3, our model comprises three components: a visual feature extractor, a text tokenizer, and an LLM decoder.
Visual Feature Extractor. The visual feature extractor adopts a dual-branch design: a spatial branch with image encoder $\mathcal{F}_{I}$ (i.e., SigLIP) processes key frames, while a temporal branch with pre-trained motion encoder $\mathcal{F}_{M}$ (i.e., SlowFast) analyzes frame sequences. Both branches employ dedicated projection layers $\mathcal{P_{I}}$ and $\mathcal{P_{F}}$ (i.e., two-layer MLPs) to map spatial and temporal features into visual tokens aligned with language space. Specifically, given an input video $\bm{x}=\{\bm{x}_{i}\}_{i=0}^{N-1}$ containing $N$ frames at frame rate $r$, we first partition it into $N_{c}=\lfloor N/r\rfloor$ continuous chunks $\{\bm{c}_{k}\}^{N_{c}-1}_{k=0}$, where each chunk $\bm{c}_{k}=\{x_{j}\}^{(k+1)*r}_{j=k*r}$ spans $r$ frames. Spatial features $\bm{f}^{s}_{k}$ are extracted from the first frame $\bm{x}_{kr}$ of each chunk, while temporal features $\bm{f}^{t}_{k}$ are computed over all frames in $c_{k}$. The feature extraction process is formally expressed as:
$\bm{f}^{s}_{k}$ $=\mathcal{P}_{I}(\mathcal{F}_{I}(\bm{x}_{kr})),\quad\bm{f}^{t}_{k}=\mathcal{P}_{M}(\mathcal{F}_{M}(\bm{c}_{k})),$ (1)
$\bm{f}^{v}$ $=\mathrm{Concat}\left([{\bm{f}^{s}_{k}},{\bm{f}^{t}_{k}}]_{k=0}^{N_{c}-1}\right),$
where $\bm{f}^{v}$ is the extracted visual features of $\bm{x}$. Given a video pair $(\bm{x}^{A},\bm{x}^{B})$, we can derive the visual features $(\bm{f}^{v}_{A},\bm{f}^{v}_{B})$.
Feature Fusion via the LLM. Given an input prompt $\bm{p}$, we first encode it into text tokens $\bm{f}^{p}=\mathcal{T}(\bm{p})$ using tokenizer $\mathcal{T}$. The visual features of a video pair $(\bm{f}^{v}_{A},\bm{f}^{v}_{B})$ are then concatenated with $\bm{f}^{t}$ and fed to a pretrained LLM decoder (i.e., Qwen-2) for multimodal fusion to derive the output response for quality ranking:
$\bm{r}$ $=\mathcal{L}(\bm{f}^{v}_{A},\bm{f}^{v}_{B},\bm{f}^{p}),$ (2)
where $\bm{r}$ is expected to belong to {“better”, “worse”} for hard ranking and {“superior”, “better”, “similar”, “worse”, “inferior”} for soft ranking.
Table 2: Performance comparison of our models against competitive opinion-unaware and supervised methods in the zero-shot setting. The best results are highlight in boldface, the second best is underlined. NA in the table indicates unavailable results. “Overall” represents the weighted average results based on the number of videos in each dataset. PLVD-P1/P2/P3 represent PLVD-Part1/ Part2/ Part3 respectively, A ✔ in the ”Label” column indicates that human-labeled data is used for training
[…]
# of videos 7,182 3,573 1,200 585 1,020 -
Methods Training data Label SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
Traning-free / Opinion-unaware VQA Methods
NIQE [35] None 0.442 0.332 0.489 0.459 0.541 0.553 0.596 0.628 0.278 0.290 0.457 0.395
IL-NIQE [67] 0.483 0.362 0.418 0.397 0.512 0.530 0.484 0.532 0.291 0.323 0.454 0.390
VIIDEO [36] 0.080 0.080 0.009 0.019 0.299 0.300 0.033 0.215 0.058 0.154 0.077 0.095
STEM [17] 0.206 0.243 0.434 0.381 0.619 0.627 0.594 0.629 0.284 0.318 0.325 0.336
TPQI [24] NA NA NA NA 0.556 0.549 0.636 0.645 0.111 0.218 0.411 0.449
BUONA-VISTA [59] NA NA NA NA 0.760 0.760 0.784 0.794 0.525 0.556 0.680 0.693
[…]
MinimalisticVQA(VII) [49] LSVQ [65] 0.861 0.859 0.740 0.784 0.843 0.841 0.757 0.813 0.775 0.779 0.817 0.830
MinimalisticVQA(IX) [49] LSVQ [65] 0.885 0.882 0.792 0.828 0.862 0.859 0.775 0.821 0.826 0.821 0.849 0.859
FAST-VQA [57] LSVQ [65] 0.880 0.880 0.781 0.813 0.859 0.854 0.826 0.845 0.730 0.747 0.838 0.849
DOVER [60] LSVQ [65] 0.878 0.866 0.782 0.813 0.874 0.869 0.817 0.840 0.771 0.781 0.842 0.845
Q-Align [61] fused [14, 8, 25, 65, 37] 0.886 0.884 0.761 0.822 0.876 0.878 0.783 0.819 0.834 0.846 0.844 0.861
[…]
Hard Ranking PLVD-P1 (500k) 0.883 0.866 0.799 0.817 0.886 0.869 0.791 0.820 0.839 0.844 0.854 0.850
Soft Ranking (Stage 1) PLVD-P1 (500k) 0.886 0.880 0.803 0.830 0.891 0.888 0.797 0.832 0.845 0.849 0.858 0.863
Soft Ranking (Stage 2) PLVD-P1/ P2 (600k) 0.887 0.880 0.802 0.830 0.893 0.892 0.798 0.833 0.856 0.859 0.859 0.864
Soft Ranking (Stage 3) PLVD-P1/ P2/ P3 (700k) 0.888 0.884 0.806 0.835 0.894 0.893 0.801 0.836 0.854 0.861 0.861 0.868
Out of Distribution Datasets LIVE-YT-Gaming CGVDS LIVE-YT-HFR Waterloo-IVC-4K KVQ Overall
# of videos 600 357 480 1,200 2,926 -
Methods Training data Label SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
Traning-free / Opinion-unaware VQA Methods
NIQE [35] None 0.240 0.247 0.473 0.496 0.354 0.413 0.048 0.002 0.163 0.114 0.183 0.154
IL-NIQE [67] 0.200 0.168 0.340 0.303 -0.081 -0.040 0.079 -0.008 0.056 0.006 0.083 0.036
VIIDEO [36] 0.077 -0.199 0.157 -0.257 0.276 0.244 0.114 0.078 0.082 0.019 0.110 0.010
STEM [17] 0.103 0.111 0.498 0.492 0.288 0.317 0.184 0.097 0.123 0.104 0.172 0.147
TPQI [24] NA NA NA NA NA NA NA NA NA NA NA NA
BUONA-VISTA [59] NA NA NA NA NA NA NA NA NA NA NA NA
[…]
MinimalisticVQA(VII) [49] LSVQ [65] 0.596 0.682 0.681 0.733 0.061 0.130 0.275 0.338 0.604 0.659 0.490 0.551
MinimalisticVQA(IX) [49] LSVQ [65] 0.686 0.746 0.797 0.816 0.301 0.388 0.459 0.502 0.615 0.661 0.574 0.622
FAST-VQA [57] LSVQ [65] 0.631 0.677 0.725 0.747 0.326 0.415 0.327 0.363 0.518 0.526 0.486 0.512
DOVER [60] LSVQ [65] [37] 0.647 0.728 0.694 0.747 0.360 0.465 0.368 0.418 0.559 0.593 0.519 0.569
Q-Align [61] fused [14, 8, 25, 65, 37] 0.611 0.681 0.756 0.798 0.329 0.342 0.414 0.497 0.613 0.655 0.555 0.606
[…]
Hard Ranking PLVD-P1 (500k) 0.705 0.742 0.794 0.801 0.496 0.550 0.492 0.542 0.640 0.670 0.613 0.648
Soft Ranking (Stage 1) PLVD-P1 (500k) 0.697 0.752 0.799 0.829 0.481 0.525 0.552 0.614 0.690 0.725 0.650 0.693
Soft Ranking (Stage 2) PLVD-P1/ P2 (600k) 0.717 0.763 0.810 0.834 0.515 0.611 0.633 0.696 0.749 0.764 0.704 0.741
Soft Ranking (Stage 3) PLVD-P1/ P2/ P3 (700k) 0.703 0.761 0.792 0.823 0.559 0.644 0.670 0.728 0.754 0.770 0.716 0.752
4.1.1 Training and Inference Pipelines
Training. LMM-PVQA is optimized via standard supervised fine-tuning (SFT). To enhance training efficiency, the LLM parameters are kept frozen to preserve their intrinsic reasoning capabilities. Similarly, the pretrained SlowFast motion encoder remains frozen due to its proven effectiveness in capturing temporal representations [49]. Trainable parameters are confined to the image encoder $\mathcal{F_{I}}$ and feature projection layers $\mathcal{P}_{I}$ and $\mathcal{P}_{M}$, thus enabling domain-specific visual feature adaptation.
Furthermore, we propose an iterative self-improvement training strategy. The training process consists of three stages. First, the base model is trained on the PLVD-Part1 dataset. Next, the trained model acts as an enhanced judge, annotating video pairs in the PLVD-Part2 dataset alongside existing judges. The combined PLVD-Part1 and PLVD-Part2 datasets are then used to train a second-stage model. This process is repeated for a third-stage model. As demonstrated in Section 5.2, this iterative strategy progressively enhances out-of-distribution performance.
Inference. Given that the model primarily predicts relative quality rankings between video pairs, a practical conversion to absolute quality scores is required. To achieve this, we adopt an adaptive soft comparison method proposed in  [71], which first computes a soft probability matrix across ranking categories by comparing the test video against anchor videos, followed by maximum a posteriori (MAP) [51] estimation under Thurstone’s Case V model [50] to obtain calibrated quality scores.
For anchor selection, the range of quality scores of PLVD-VQA is partitioned into five quality intervals. Within each interval, the video exhibiting the lowest inter-model score variance is selected as the anchor, ensuring stable reference points for score derivation. We list the detailed calculation procedure in the supplement file.
5 Experiments
5.1 Experimental Setups
Implementation Details For each stage of training, we initialize the image encoder, image projector, and LLM with LLaVA-ov-chat (7B) weights [21] and the motion encoder with SlowFast pre-trained weights. The model is trained for one epoch on four NVIDIA A800 GPUs with a total batch size of $4$ and a learning rate of $1$e-$5$. Inference is performed on two NVIDIA RTX 3090 GPUs.
Benchmark Datasets The proposed methods is evaluated on ten VQA benchmarks, categorized into in-domain and out-of-distribution (OOD) groups to systematically assess their generalization. The in-domain evaluation benchmarks include LSVQ Test [65], LSVQ 1080p [65], KoNViD-1k [13], LIVE-VQC [47], and YouTube-UGC [56], all of which contain UGC videos. The OOD datasets consist of LIVE-YT-Gaming [66], CGVDS [44], LIVE-YT-HFR [30], Waterloo-IVC-4K [23], and KVQ [29]. Specifically, LIVE-YT-Gaming and CGVDS contain gaming videos, which exhibit significant differences in content compared to natural UGC videos. LIVE-YT-HFR includes videos with varying frame rates, focusing on temporal-related distortions. Waterloo-IVC-4K and KVQ primarily evaluate compression artifacts, where Waterloo-IVC-4K contains compressed high-resolution videos at different spatial resolutions, while KVQ comprises compressed UGC videos processed with various video enhancement techniques. We introduce details of these datasets in the supplement file.
Evaluation Criteria We adopt two widely used criteria to evaluate the performance of VQA models: Spearman Rank Correlation (SRCC) and Pearson Linear Correlation (PLCC), which indicate the prediction monotonicity and prediction linearity, respectively.
Competing Methods We compare our model with six opinion-unaware BVQA methods, including NIQE [35], IL-NIQE [67], VIIDEO [36], STEM [17], TPQI [24], and BUONA-VISTA [59], as well as five supervised VQA methods: Minimalistic-VQA (VII) [49], Minimalistic-VQA (IX) [49], FAST-VQA [57], DOVER [60], and Q-Align [61]. These five supervised VQA models serve as strong baselines, as they are used to pseudo-label video quality and can be regarded as teacher models. The supervised models are trained on the LSVQ training set.
Table 3: Supervised performance comparison of our model against competitive methods. The experimental results on KoNViD-1k, YouTube-UGC, and LIVE-VQC are obtained using 10-fold cross-validation. “Base” denotes training LMM-PVQA without self-supervised pretraining
[…]
TLVQM [19] 0.772 / 0.774 0.589 / 0.616 0.732 / 0.724 0.670 / 0.691 0.685 / 0.692
VIDEAL [52] 0.794 / 0.783 0.545 / 0.554 0.751 / 0.741 0.630 /0.640 0.687 / 0.709
VSFA [22] 0.801 / 0.796 0.675 / 0.704 0.794 / 0.799 0.718 / 0.771 0.787 / 0.789
PatchVQ [65] 0.827 / 0.828 0.711 / 0.739 0.791 / 0.786 0.827 / 0.837 NA / NA
SimpleVQA [48] 0.867 / 0.861 0.764 / 0.803 0.856 / 0.860 0.845 / 0.859 0.847 / 0.856
FAST-VQA [57] 0.880 / 0.880 0.781 / 0.813 0.891 / 0.892 0.849 / 0.865 0.855 / 0.852
DOVER [60] 0.878 / 0.866 0.782 / 0.813 0.908 / 0.910 0.860 / 0.875 0.841 / 0.851
MinimalisticVQA [49] 0.881 / 0.879 0.781 / 0.820 0.889 / 0.890 0.842 / 0.854 0.890 / 0.891
Soft Ranking (Base) 0.880 / 0.864 0.790 / 0.814 0.896 / 0.877 0.840 / 0.853 0.858 / 0.848
Soft Ranking (Stage 1) 0.907 / 0.904 0.832 / 0.857 0.911 / 0.908 0.884 / 0.894 0.910 / 0.911
Table 4: Experimental results of ablation study. “Spatial” and “Motion” denote the spatial branch and the temporal branch of the visual feature extractor. In Stage 1, if PLVD-DS (Part1) is not specified, the model is trained solely on PLVD-VQA (Part1). In Stage 2, “w/o label refinement” and “w/ label refinement” indicate whether the trained model is used to label the data in PLVD-Part2
[…]
Spatial Motion PLVD-DS w/o label refinement w/ label refinement SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
0.872 0.862 0.787 0.814 0.868 0.864 0.783 0.819 0.834 0.839 0.843 0.846
0.883 0.872 0.804 0.827 0.883 0.876 0.799 0.830 0.842 0.843 0.855 0.857
0.886 0.880 0.803 0.830 0.891 0.888 0.797 0.832 0.845 0.849 0.858 0.863
[…]
0.887 0.880 0.802 0.830 0.893 0.892 0.798 0.833 0.856 0.859 0.859 0.864
Stage 1 Stage 2 LIVE-YT-Gaming CGVDS LIVE-YT-HFR Waterloo-IVC-4K KVQ Overall
Spatial Motion PLVD-DS w/o label refinement w/ label refinement SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
0.678 0.735 0.754 0.800 0.334 0.413 0.404 0.459 0.616 0.656 0.561 0.610
0.688 0.736 0.769 0.808 0.446 0.477 0.425 0.470 0.619 0.662 0.579 0.622
0.697 0.752 0.799 0.829 0.481 0.525 0.552 0.614 0.690 0.725 0.650 0.693
[…]
0.717 0.763 0.810 0.834 0.515 0.611 0.633 0.696 0.749 0.764 0.704 0.741
5.2 Performance Analysis
The performance of all compared methods and our proposed models is summarized in Table 4.1, with detailed analysis conducted from the following perspectives:
Hard Ranking v.s. Soft Ranking. The model trained with soft ranking marginally outperforms its hard-ranking counterpart on in-domain VQA benchmarks and CGVDS. Moreover, it achieves significantly better performance on Waterloo-IVC-4K and KVQ, while exhibiting slightly inferior results on LIVE-YT-Gaming and LIVE-YT-HFR. Overall, the soft ranking strategy incorporates “quality distance” information into the training process, enabling the model to easily learn quality comparison capabilities and achieve better performance.
Iterative Self-Improvement Training Strategy. Consistent performance improvements are observed across both in-domain and OOD datasets as the model advances through training stages. This empirical evidence demonstrates that our iterative training strategy effectively enhances model capability through progressive self-learning. Notably, substantial gains are achieved on challenging VQA benchmarks where existing models struggle: the LIVE-YT-HFR, Waterloo-IVC-4K, and KVQ datasets exhibit relative improvements of $16.2\%$, $21.4\%$, $9.3\%$ in SRCC respectively, after three-stage refinement.
In-Domain Performance Comparison.
We observe that all proposed model variants outperform the five competing VQA models, which are used for label annotation in Section 3.2.1, confirming our self-supervised approach successfully distills ensemble knowledge from its teacher models on unlabeled data. However, our models are slightly outperformed by FAST-VQA and DOVER on LIVE-VQC, which exhibits more complex temporal distortions compared to other datasets [49]. We attribute this performance gap to architectural differences: FAST-VQA and DOVER fine-tune a 3D DNN (i.e., VideoSwin) specifically optimized for spatiotemporal feature extraction, whereas our framework handles temporal information using a pre-trained SlowFast model.
OOD Performance Comparison. Our models significantly outperform competing approaches on the OOD benchmarks, achieving a $24.7\%$ improvement in SRCC for overall OOD evaluation compared to the competing approaches. We attribute the strong OOD performance to two key factors. First, our training datasets include compression distortions (H.264 and H.265), and the synthetic quality ranking labels provide useful supervision for assessing compression-related video quality for these datasets. Second, our iterative self-improvement training strategy enables progressive adaptation to unseen distortion types, such as frame rate inconsistencies in LIVE-YT-HFR, AVS/VP9 compression artifacts in Waterloo-IVC-4K, and enhancement-induced distortions in KVQ—none of which are present in the training data. These results demonstrate the effectiveness of the proposed self-supervised VQA framework in real-world quality assessment scenarios.
5.3 Supervised Performance Comparison
We conduct supervised fine-tuning of our self-supervised trained model on the in-domain VQA benchmarks, as shown in Table 3. The results reveal that: (1) Our fine-tuned model surpasses all competing VQA baselines with 3.5% higher SRCC, and (2) it achieves $6.4\%$ improvement over the baseline trained without self-supervised initialization. These results confirm that self-supervised representation learning critically enhances downstream fine-tuning effectiveness.
5.4 Ablation Study
We conduct the ablation experiments to validate the effectiveness of LMM-PVQA, with results shown in Table 4.
Motion Encoder. Incorporating the motion encoder for temporal distortion representation yields consistent improvements across all benchmarks, particularly achieving $33.5\%$ SRCC improvement on LIVE-YT-HFR—the benchmark specifically designed for high frame-rate distortion analysis. This verifies that explicit motion modeling is essential for capturing temporal-related distortions.
Synthetic Distortion Data. The incorporation of the PLVD-DS dataset also demonstrates consistent performance improvements across all benchmarks, yielding marginal gains on in-domain benchmarks while achieving substantial enhancements on OOD benchmarks. It can be attributed that the PLVD-VQA dataset have already equipped the model with quality assessment capabilities for in-domain content, whereas the PLVD-DS dataset help mitigate the domain gap on OOD benchmarks by simulating common artificial degradations (e.g., compression).
Iterative Self-Improvement Training Strategy. Since our iterative self-improvement training strategy introduces new training samples, it is critical to verify that performance gains stemming from the strategy itself rather than additional data. To isolate this effect, we train a Stage 2 model using the same base training set (PLVD-Part1/Part2), where videos in PLVD-Part2 are labeled by five VQA judges, without incorporating the Stage 1 model as the new judge. Experimental results confirm that merely increasing training data (without iterative refinement) fails to surpass the Stage 1 model’s performance, thereby validating the effectiveness of our strategy in enabling self-improvement through iterative optimization.
6 Conclusion
We propose a self-supervised learning framework for VQA to mitigate dependence on human-annotated datasets. Our approach constructs a large-scale video pair dataset labeled via established VQA models and synthetic distortion simulations, then adopts a learning-to-rank paradigm to learn quality assessment capabilities. By integrating an iterative self-improvement training strategy, our model progressively enhances its evaluation performance through self-learning. Our method achieves state-of-the-art zero-shot performance on both in-domain and OOD benchmarks, demonstrating its effectiveness.
Limitations. Our method is inherently designed for pairwise quality comparison. To obtain absolute quality scores for a single video, it must be compared against multiple anchor videos (five in this study), which significantly increases inference time.
Future Work. We plan to explore more automated video pair annotation strategies, such as leveraging expert-domain VQA models (e.g., VMAF for video compression), utilizing LMMs with carefully designed prompt engineering, and employing text-to-video generation algorithms to synthesize videos of varying quality through specified prompts, thus substantially expanding the diversity and scale of training data. Furthermore, we intend to extend our framework to incorporate additional modalities, such as images and audio, to develop a more generalizable quality assessment model.