### v4

# Breaking Annotation Barriers: Generalized Video Quality Assessment via Ranking-based Self-Supervision

Abstract
Video quality assessment (VQA) is essential for quantifying perceptual quality in various video processing workflows, spanning from camera capture systems to over-the-top streaming platforms. While recent supervised VQA models have made substantial progress, the reliance on manually annotated datasets—a process that is labor-intensive, costly, and difficult to scale up—has hindered further optimization of their generalization to unseen video content and distortions. To bridge this gap, we introduce a self-supervised learning framework for VQA to learn quality assessment capabilities from large-scale, unlabeled web videos. Our approach leverages a learning-to-rank paradigm to train a large multimodal model (LMM) on video pairs automatically labeled via two manners, including quality pseudo-labeling by existing VQA models and relative quality ranking based on synthetic distortion simulations. Furthermore, we introduce a novel iterative self-improvement training strategy, where the trained model acts an improved annotator to iteratively refine the annotation quality of training data. By training on a dataset $10\times$ larger than the existing VQA benchmarks, our model: (1) achieves zero-shot performance on in-domain VQA benchmarks that matches or surpasses supervised models; (2) demonstrates superior out-of-distribution (OOD) generalization across diverse video content and distortions; and (3) sets a new state-of-the-art when fine-tuned on human-labeled datasets. Extensive experimental results validate the effectiveness of our self-supervised approach in training generalized VQA models. The datasets and code will be publicly released to facilitate future research. Our code and database will be available at https://github.com/clh124/LMM-PVQA.
Figure 1: Overview of our work. (a) Existing state-of-the-art VQA models exhibit poor out-of-distribution performance. (b) We construct a large-scale VQA dataset consisting of 700700k video pairs, sampled from multiple social media platforms, covering more than 20 content categories. (c) We explore two strategies for automatically annotating the relative quality of video pairs. (d) Our proposed model undergoes iterative training to enhance generalization performance.
footnotetext: *Equal contribution. Project lead.
1 Introduction
Video quality assessment (VQA) [35] This work focuses on no-reference (NR) or blind VQA, which assesses video quality without relying on additional reference information. plays an important role in modern video processing systems, delivering objective quality measurements used to optimize end-user Quality of Experience (QoE). With the advances in deep neural networks (DNNs) [15, 10, 29] and the increasing availability of human-annotated VQA datasets [16, 50, 60, 70], current VQA models [52, 61, 64, 65] have achieved significant progress through supervised learning. Nevertheless, supervised learning inherently faces a limitation: the generalization of the VQA models heavily depends on the diversity of the training data. For example, even top-tier VQA models [52, 61, 64, 65] exhibit significant performance drops in out-of-distribution evaluations, as illustrated in Fig. 1(a).
Therefore, the development of specialized VQA datasets remains critical to address the growing diversity of emerging media formats and their associated distortions [76, 75, 49, 67, 26, 33, 71]. However, constructing such datasets is highly resource-intensive. A standardized subjective experiment comprises two key phases: test sample curation and subjective quality annotation. The test sample curation phase necessitates rigorous selection of representative video samples, as inadequate sampling strategies risk producing oversimplified datasets (i.e.,easy dataset” problem [52, 4]) and may induce model overfitting. Meanwhile, subjective annotation—though vital—is laborious and costly. International Telecommunication Union (ITU) standards (e.g., ITU-T P.910 [18]) outline specific recommendations for experimental setups, including display conditions, stimulus duration, subject count, and rating methodologies. These constraints, though necessary for statistically meaningful results, impede large-scale dataset expansion due to prohibitive annotation costs.
Self-supervised or unsupervised learning [13], which eliminates the need for human-labeled annotations, is a potential solution in VQA to mitigate the high costs of subjective experiments. However, current self-supervised VQA methods [7, 6, 8, 34, 36] still lag behind their supervised counterparts in performance. Typical implementations of current self-supervised VQA methods [6, 34] employ contrastive learning frameworks with proxy tasks such as distortion type/severity classification on synthetically generated data. However, it suffers from two shortcomings: (1) they fail to capture visual content and aesthetic characteristics relevant to perceptual quality assessment, and (2) they inadequately model real-world authentic distortion patterns that follow complex nonlinear degradation processes.
In this paper, we propose a novel self-supervised learning framework for VQA to address the aforementioned challenges. Specifically, we reformulate quality regression as a ranking problem, enabling the model to learn quality assessment capabilities through pairwise comparisons. This approach is motivated by the observation that pairwise ranking is more intuitive and reliable than absolute quality ratings for human evaluators. We then explore two strategies to automatically label the relative quality of video pairs. The first leverages existing VQA model as the “judge” to determine the relative quality, where we ensemble the results from multiple judges to mitigate inherent evaluation biases in individual models. This process is iterative: once a new VQA model is trained, it can serve as an enhanced judge, achieving self-improvement through repeated refinement. The second approach relies on synthetic distortion simulations, where we introduce various types of distortions at different severity levels and utilize these severity levels to establish relative quality rankings.
We adopt a large multimodal model (LMM) integrated with a fixed motion encoder and a learnable motion projector as our VQA model. Given two videos and a text prompt as input, the model outputs a textual judgment indicating whether the first video has higher or lower quality than the second. We employ a standard supervised fine-tuning (SFT) strategy to train the proposed VQA model using automatically annotated video pairs. Training is conducted in three iterative stages with progressively increasing sample sizes of 500500k, 600600k, and 700700k video pairs, respectively. We evaluate the proposed model on ten diverse VQA benchmarks. Experimental results show that the model achieves zero-shot performance comparable to, or even surpassing, existing supervised VQA models, while demonstrating superior out-of-distribution generalization. Moreover, after fine-tuning on human-annotated datasets, the model significantly outperforms state-of-the-art supervised approaches.
Our key contributions are summarized as follows:
  • We construct a large-scale VQA dataset comprising 700k video pairs with both authentic and synthetic distortions, each annotated with automatically generated quality labels.
  • We propose a self-supervised VQA framework that enables the model to learn quality assessment capabilities through pairwise comparisons, enhanced by an iterative training strategy for continuous self-improvement.
  • Our model achieves strong zero-shot performance across multiple VQA benchmarks, highlighting its effectiveness and generalization in video quality assessment.
2 Related Work
2.1 VQA Models
Supervised VQA. Early supervised VQA models, such as V-BLIINDS [46], TLVQM [22], and VIDEVAL [55], are typically knowledge-driven. These methods extract quality-related features based on natural scene statistics (NSS) [37], blurriness [22], motion vectors [21], optical flow [3], and other handcrafted cues, followed by training a machine learning-based regressor, such as a support vector regressor (SVR) or random forest regressor, to derive quality scores.
With the popularity of DNNs, some VQA methods, such as VSFA [25], Li22 [23], PatchVQA [70], etc., employ pre-trained DNNs as feature extractors to derive quality representations from video frames, followed by training quality regressors to map the extracted features into quality scores. Commonly used feature extractors include image classification models [15], image quality assessment (IQA) models [73, 69], and action recognition models [12], while commonly used quality regressors often consist of GRUs [25], Transformers [62], and InceptionTime [70], etc.
Many recent works explore the use of 2D or 3D convolutional neural networks (CNNs) [15], Vision Transformers (ViTs) [10], and LMMs [65] as feature extractors, fine-tuning them in an end-to-end manner to achieve state-of-the-art performance. For instance, SimpleVQA [51] and MinimalisticVQA [52] perform end-to-end training of the spatial feature extractor (e.g., ResNet [15], Swin [29]) while adopting a temporal extractor (i.e., SlowFast [12]) with fixed pre-trained weights. FAST-VQA [61] trains a 3D DNN (i.e., VideoSwin [31]) in an end-to-end fashion, and DOVER [64] further extends FAST-VQA with an aesthetic evaluation branch based on a 2D CNN (i.e., ConvNet [30]). Q-Align [65] fine-tunes the visual encoder of an LMM with a predefined text-based quality inquiry prompt.
Unsupervised and self-supervised VQA. A class of popular unsupervised or self-supervised VQA approaches [6, 34, 36, 66] aims to learn quality-aware feature representations from scratch, followed by fine-tuning a linear projector with human-annotated labels to serve as a quality score regressor. For example, CSPT [6] and CONVIQT [34] adopt contrastive learning frameworks with proxy tasks such as next-frame feature discrimination and distortion type/severity classification to learn the quality-aware representation. QPT V2 [66], on the other hand, employs an encoder-decoder architecture to reconstruct pristine videos from distorted ones, thereby learning distortion-aware representations.
Another line of research focuses on developing opinion-unaware VQA methods, which aligns with the objective of this work. Some knowledge-driven VQA methods [38, 72, 27, 39, 20] estimate video quality by directly measuring the distribution distance of specific features between distorted and pristine videos. For instance, NIQE [38] assesses the spatial quality of a test image by computing the distance between the Multivariate Gaussian model fitted to local features of the test image and pristine natural images. TPQI [27] captures temporal distortions by analyzing the straightness and compactness of video trajectories within perceptual domains of the human visual system. Recent studies [64, 1] leverage visual-language models to achieve zero-shot I/VQA. For example, BUONA-VISTA [63] employs CLIP [44] to estimate the relatively probabilities of text promt “high quality” compared to “low quality”, and combines these with NIQE and TPQI scores to assess video quality.
2.2 VQA Datasets
The progress of data-driven VQA models heavily relies on VQA datasets annotated by human subjects. Early VQA datasets, such as LIVE-VQA [48] and CSIQ-VQA [57], primarily focus on compression and transmission distortions in professionally generated content (PGC). However, these datasets contain a limited number of source videos and only a few hundred distorted samples. While they have contributed to the development of knowledge-driven VQA methods, their scale and diversity make them inadequate for training data-driven VQA models. With the rise of social media platforms, the focus has shifted to user generated content (UGC) videos characterized by natural and in-the-wild distortions. Current mainstream UGC VQA datasets, such as KoNVid-1k [16], YouTube-UGC [60], and LSVQ [70], contain thousands to tens of thousands of videos, providing large-scale benchmarks that have significantly advanced research in data-driven VQA. Meanwhile, with the emergence of new media formats, specialized VQA datasets have been developed to address domain-specific challenges, such as gaming videos [71], 360-degree videos [67], 4K videos  [26], high-frame-rate videos [33], AIGC videos [75], etc., each designed to facilitate the development of expert VQA models tailored to specific video types.
3 Pairwise-Labeled Video Dataset
This section describes the construction of the large-scale pairwise-labeled video dataset (PLVD) in detail.
Figure 2: Distribution of nine metrics across mainstream UGC datasets (LSVQ, KoNVid-1k, YouTube-UGC, LIVE-VQC), as well as our dataset before and after sampling.
3.1 Video Collection
Source Video Collection. We create a large-scale dataset comprising $3$ million videos collected from popular social media platforms, including YouTube, TikTok, Youku, and Bilibili. This dataset encompasses a wide range of content categories and distortion scenarios, such as vlogs, gaming videos, animations, live streams, etc., ensuring a representative and diverse collection with varying quality conditions. We provide the detail analysis of source videos in the supplement file.
Candidate Video Sampling. We select a subset of videos from the collected pool using a mixed-integer programming approach [56] to match target distributions defined by nine low-level metrics that quantify visual characteristics of datasets, including blockiness [45], blur [41], contrast [43], noise, flickering [42], colorfulness [14], luminance, temporal information (TI) [19], and spatial information (SI) [19]. We provide the calculation of these metrics in the supplement file. Our target distribution mirrors the aggregated distributions of mainstreaming UGC datasets—LSVQ [70], KonViD-1k [16], YouTube-UGC [60], and LIVE-VQC [50]—to ensure compatibility with existing benchmarks. Finally, we sample 438438k videos to enhance diversity in content and scenes while maintaining close alignment with mainstream UGC datasets. As illustrated in Fig. 2, the distributions of the sampled subset exhibit strong consistency with that of public VQA datasets across all nine metrics.
3.2 Pairwise Quality Annotation
We explore two strategies to automatically annotate the relative quality of video pairs: (1) quality pseudo-labeling using existing VQA models, and (2) relative quality ranking based on synthetic distortion simulations. We define two types of ranking labels: hard ranking and soft ranking. Specifically, hard ranking consists of two categories, “better” and “worse”, while soft ranking provides finer-grained comparisons with five levels: “superior”, “better”, “similar”, “worse”, and “inferior”, to reflect varying degrees of relative quality.
Table 1: Statistics of raw videos and video pairs in the PLVD dataset. The values in each cell indicate the number of videos or video pairs in PLVD-Part1/-Part2/-Part3
Videos Video Pairs
PLVD-VQA 200k/100k/50k 250k/85k/85k
PLVD-DS Spatial 50k/2k/2k 160k/5k/5k
Temporal 20k/1k/1k 40k/5k/5k
Compression 10k/1k/1k 50k/5k/5k
Total 280k/384k/438k 500k/600k/700k
3.2.1 Pseudo-labeling using VQA Models
Inspired by recent LLM-as-a-Judge methods [58] that leverage large language models (LLM) to model user preference for LLM alignment, we propose utilizing established VQA models as judges to evaluate the relative quality of video pairs. We refer to this subset of data as PLVD-VQA. Specifically, we choose five state-of-the-art VQA models: Minimalistic-VQA (VII) [52], Minimalistic-VQA (IX) [52], FAST-VQA [61], DOVER [64], and Q-Align [65], all trained on LSVQ [70], as our initial judges. For a video pair $(x^{A},x^{B})$, each judge model $\mathcal{J}_{i}$ generates quality scores $j^{A}_{i}$ and $j^{B}_{i}$. For hard ranking, the judgment is determined by comparing the mean quality scores $\overline{j}^{A}$ and $\overline{j}^{B}$. If $\overline{j}^{A}>\overline{j}^{B}$, we annotate $x^{A}$ as higher quality than $x^{B}$ (label: “better”); otherwise, $x^{A}$ is labeled as worse quality (label: “worse”).
For soft ranking, we further compute the score variance for the video pair $\sigma^{2}_{A}$ and $\sigma^{2}_{B}$. Assuming the quality difference $\Delta=\overline{j}^{A}-\overline{j}^{B}$ follows a Gaussian distribution $\mathcal{N}(\Delta;0,\sigma^{2}_{\Delta})$, where $\sigma_{\Delta}=\sqrt{\sigma^{2}_{A}+\sigma^{2}_{B}}$, we assign labels based on statistical significance thresholds adapted from [77]: pairs are labeled as “superior” if $\Delta>2\sigma_{\Delta}$, “better” if $\sigma_{\Delta}<\Delta\leq 2\sigma_{\Delta}$, “similar” if $\sigma_{\Delta}<\Delta\leq\sigma_{\Delta}$, “worse” if $2\sigma_{\Delta}<\Delta\leq-\sigma_{\Delta}$, and “inferior” if $\Delta\leq-2\sigma_{\Delta}$.
3.2.2 Quality Ranking via Distortion Simulations
Figure 3: Overall framework of the proposed LMM-PVQA model. Compared to the standard LMM, our visual feature extractor adopts a dual-branch design with an additional motion encoder and projector to capture temporal-related distortions. The model takes a video pair and a text prompt as inputs and generates a text output specifying which video has better quality.
We introduce synthetic distortions to simulate typical degradation that may occur in real-world scenarios, which are categorized into three types: spatial distortions, temporal distortions, and streaming distortions. Specifically, spatial distortions include resolution resizing, Gaussian blur, Gaussian noise, darkening, and brightening, which simulate capture-related artifacts. Temporal distortions consist of jitter and stuttering, mimicking playback issues commonly observed in practical settings. Streaming distortions involve H.264 and H.265 compression, reflecting compression artifacts introduced by streaming media platforms. We denote this subset of data as PLVD-DS. The detailed generation of synthetic distortions are provided in the supplement file.
We leverage distortion severity levels (e.g., constant rate factor for compression) as pseudo-labels to infer relative quality. Given a primary video $x^{0}$ and a synthetic distortion simulator $\mathcal{S}$, we degrade $x^{0}$ across $N_{\mathcal{S}}$ severity levels to generate distorted videos $\{x_{\mathcal{S}}^{i}\}^{N_{\mathcal{S}}}_{i=1}$. Pairs $(x_{\mathcal{S}}^{i},x_{\mathcal{S}}^{j})$ are randomly sampled. For hard ranking, pairs are directly annotated as “better” if $i<j$ or “worse” otherwise. For soft ranking, pairs with a severity difference |ij|>1|i-j|>1 are labeled as “superior” or “inferior” depending on the relative order of $i$ and $j$, while pairs with |ij|=1|i-j|=1 receive “better” or “worse”. The “similar” label is intentionally excluded, as ij=0i-j=0 implies identical videos.
3.2.3 Label Refinement
The label quality of the PLVD-VQA dataset inherently depends on the performance of judges. To address this dependency, we introduce an iterative label refinement framework. Specifically, once a new VQA model is trained, we treat it as an improved judge $\mathcal{J}^{\prime}$ and reapply the annotation pipeline of PLVD-VQA to iteratively refine video pair labels.
In summary, the PLVD dataset comprises 700700k annotated video pairs, partitioned into three subsets of 500500k, 100100k, and 100100k pairs, denoted as PLVD-Part1, PLVD-Part2, PLVD-Part3. The latter two subsets are dedicated to iterative label refinement via the trained model. A detailed breakdown of PLVD, including pair types and the corresponding number of videos, is provided in Table 1.
4 Proposed Method
We aim to train a VQA model on an unlabeled video dataset to compute the perceptual quality score of a video. To achieve this goal, we reformulate the VQA regression task as a classification problem that distinguishes the relative quality between pairs of videos.
4.1 Model Structure
We introduce an LMM-based VQA framework for pairwise quality ranking (LMM-PVQA). As illustrated in Fig. 3, our model comprises three components: a visual feature extractor, a text tokenizer, and an LLM decoder.
Visual Feature Extractor. The visual feature extractor adopts a dual-branch design: a spatial branch with image encoder $\mathcal{F}_{I}$ (i.e., SigLIP) processes key frames, while a temporal branch with pre-trained motion encoder $\mathcal{F}_{M}$ (i.e., SlowFast) analyzes frame sequences. Both branches employ dedicated projection layers $\mathcal{P_{I}}$ and $\mathcal{P_{F}}$ (i.e., two-layer MLPs) to map spatial and temporal features into visual tokens aligned with language space. Specifically, given an input video $\bm{x}=\{\bm{x}_{i}\}_{i=0}^{N-1}$ containing $N$ frames at frame rate $r$, we first partition it into $N_{c}=\lfloor N/r\rfloor$ continuous chunks $\{\bm{c}_{k}\}^{N_{c}-1}_{k=0}$, where each chunk $\bm{c}_{k}=\{x_{j}\}^{(k+1)*r}_{j=k*r}$ spans $r$ frames. Spatial features $\bm{f}^{s}_{k}$ are extracted from the first frame $\bm{x}_{kr}$ of each chunk, while temporal features $\bm{f}^{t}_{k}$ are computed over all frames in $c_{k}$. The feature extraction process is formally expressed as:
$\bm{f}^{s}_{k}$ $=\mathcal{P}_{I}(\mathcal{F}_{I}(\bm{x}_{kr})),\quad\bm{f}^{t}_{k}=\mathcal{P}_{M}(\mathcal{F}_{M}(\bm{c}_{k})),$ (1)
$\bm{f}^{v}$ $=\mathrm{Concat}\left([{\bm{f}^{s}_{k}},{\bm{f}^{t}_{k}}]_{k=0}^{N_{c}-1}\right),$
where $\bm{f}^{v}$ is the extracted visual features of $\bm{x}$. Given a video pair $(\bm{x}^{A},\bm{x}^{B})$, we can derive the visual features $(\bm{f}^{v}_{A},\bm{f}^{v}_{B})$.
Feature Fusion via the LLM. Given an input prompt $\bm{p}$, we first encode it into text tokens $\bm{f}^{p}=\mathcal{T}(\bm{p})$ using tokenizer $\mathcal{T}$. The visual features of a video pair $(\bm{f}^{v}_{A},\bm{f}^{v}_{B})$ are then concatenated with $\bm{f}^{t}$ and fed to a pretrained LLM decoder (i.e., Qwen-2) for multimodal fusion to derive the output response for quality ranking:
$\bm{r}$ $=\mathcal{L}(\bm{f}^{v}_{A},\bm{f}^{v}_{B},\bm{f}^{p}),$ (2)
where $\bm{r}$ is expected to belong to {“better”, “worse”} for hard ranking and {“superior”, “better”, “similar”, “worse”, “inferior”} for soft ranking.
Table 2: Performance comparison of our models against competitive opinion-unaware and supervised methods in the zero-shot setting. The best results are highlight in boldface, the second best is underlined. NA in the table indicates unavailable results. “Overall” represents the weighted average results based on the number of videos in each dataset. PLVD-P1/P2/P3 represent PLVD-Part1/ Part2/ Part3 respectively, A ✔ in the ”Label” column indicates that human-labeled data is used for training
In-domain Datasets $_{\text{test}} LSVQ$$_{\text{1080p}}$ KoNViD-1k LIVE-VQC YouTube-UGC Overall
# of videos 7,182 3,573 1,200 585 1,020 -
Methods Training data Label SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
Traning-free / Opinion-unaware VQA Methods
NIQE [38] None 0.442 0.332 0.489 0.459 0.541 0.553 0.596 0.628 0.278 0.290 0.457 0.395
IL-NIQE [72] 0.483 0.362 0.418 0.397 0.512 0.530 0.484 0.532 0.291 0.323 0.454 0.390
VIIDEO [39] 0.080 0.080 0.009 0.019 0.299 0.300 0.033 0.215 0.058 0.154 0.077 0.095
STEM [20] 0.206 0.243 0.434 0.381 0.619 0.627 0.594 0.629 0.284 0.318 0.325 0.336
TPQI [27] NA NA NA NA 0.556 0.549 0.636 0.645 0.111 0.218 0.411 0.449
BUONA-VISTA [63] NA NA NA NA 0.760 0.760 0.784 0.794 0.525 0.556 0.680 0.693
Supervised VQA Methods
MinimalisticVQA(VII) [52] LSVQ [70] 0.861 0.859 0.740 0.784 0.843 0.841 0.757 0.813 0.775 0.779 0.817 0.830
MinimalisticVQA(IX) [52] LSVQ [70] 0.885 0.882 0.792 0.828 0.862 0.859 0.775 0.821 0.826 0.821 0.849 0.859
FAST-VQA [61] LSVQ [70] 0.880 0.880 0.781 0.813 0.859 0.854 0.826 0.845 0.730 0.747 0.838 0.849
DOVER [64] LSVQ [70] 0.878 0.866 0.782 0.813 0.874 0.869 0.817 0.840 0.771 0.781 0.842 0.845
Q-Align [65] fused [17, 11, 28, 70, 40] 0.886 0.884 0.761 0.822 0.876 0.878 0.783 0.819 0.834 0.846 0.844 0.861
Our Self-supervised VQA Methods: LMM-PVQA
Hard Ranking PLVD-P1 (500k) 0.883 0.866 0.799 0.817 0.886 0.869 0.791 0.820 0.839 0.844 0.854 0.850
Soft Ranking (Stage 1) PLVD-P1 (500k) 0.886 0.880 0.803 0.830 0.891 0.888 0.797 0.832 0.845 0.849 0.858 0.863
Soft Ranking (Stage 2) PLVD-P1/ P2 (600k) 0.887 0.880 0.802 0.830 0.893 0.892 0.798 0.833 0.856 0.859 0.859 0.864
Soft Ranking (Stage 3) PLVD-P1/ P2/ P3 (700k) 0.888 0.884 0.806 0.835 0.894 0.893 0.801 0.836 0.854 0.861 0.861 0.868
Out of Distribution Datasets LIVE-YT-Gaming CGVDS LIVE-YT-HFR Waterloo-IVC-4K KVQ Overall
# of videos 600 357 480 1,200 2,926 -
Methods Training data Label SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
Traning-free / Opinion-unaware VQA Methods
NIQE [38] None 0.240 0.247 0.473 0.496 0.354 0.413 0.048 0.002 0.163 0.114 0.183 0.154
IL-NIQE [72] 0.200 0.168 0.340 0.303 -0.081 -0.040 0.079 -0.008 0.056 0.006 0.083 0.036
VIIDEO [39] 0.077 -0.199 0.157 -0.257 0.276 0.244 0.114 0.078 0.082 0.019 0.110 0.010
STEM [20] 0.103 0.111 0.498 0.492 0.288 0.317 0.184 0.097 0.123 0.104 0.172 0.147
TPQI [27] NA NA NA NA NA NA NA NA NA NA NA NA
BUONA-VISTA [63] NA NA NA NA NA NA NA NA NA NA NA NA
Supervised VQA Methods
MinimalisticVQA(VII) [52] LSVQ [70] 0.596 0.682 0.681 0.733 0.061 0.130 0.275 0.338 0.604 0.659 0.490 0.551
MinimalisticVQA(IX) [52] LSVQ [70] 0.686 0.746 0.797 0.816 0.301 0.388 0.459 0.502 0.615 0.661 0.574 0.622
FAST-VQA [61] LSVQ [70] 0.631 0.677 0.725 0.747 0.326 0.415 0.327 0.363 0.518 0.526 0.486 0.512
DOVER [64] LSVQ [70] [40] 0.647 0.728 0.694 0.747 0.360 0.465 0.368 0.418 0.559 0.593 0.519 0.569
Q-Align [65] fused [17, 11, 28, 70, 40] 0.611 0.681 0.756 0.798 0.329 0.342 0.414 0.497 0.613 0.655 0.555 0.606
Our Self-supervised VQA Methods: LMM-PVQA
Hard Ranking PLVD-P1 (500k) 0.705 0.742 0.794 0.801 0.496 0.550 0.492 0.542 0.640 0.670 0.613 0.648
Soft Ranking (Stage 1) PLVD-P1 (500k) 0.697 0.752 0.799 0.829 0.481 0.525 0.552 0.614 0.690 0.725 0.650 0.693
Soft Ranking (Stage 2) PLVD-P1/ P2 (600k) 0.717 0.763 0.810 0.834 0.515 0.611 0.633 0.696 0.749 0.764 0.704 0.741
Soft Ranking (Stage 3) PLVD-P1/ P2/ P3 (700k) 0.703 0.761 0.792 0.823 0.559 0.644 0.670 0.728 0.754 0.770 0.716 0.752
4.1.1 Training and Inference Pipelines
Training. LMM-PVQA is optimized via standard supervised fine-tuning (SFT). To enhance training efficiency, the LLM parameters are kept frozen to preserve their intrinsic reasoning capabilities. Similarly, the pretrained SlowFast motion encoder remains frozen due to its proven effectiveness in capturing temporal representations [52]. Trainable parameters are confined to the image encoder $\mathcal{F_{I}}$ and feature projection layers $\mathcal{P}_{I}$ and $\mathcal{P}_{M}$, thus enabling domain-specific visual feature adaptation.
Furthermore, we propose an iterative self-improvement training strategy. The training process consists of three stages. First, the base model is trained on the PLVD-Part1 dataset. Next, the trained model acts as an enhanced judge, annotating video pairs in the PLVD-Part2 dataset alongside existing judges. The combined PLVD-Part1 and PLVD-Part2 datasets are then used to train a second-stage model. This process is repeated for a third-stage model. As demonstrated in Section 5.2, this iterative strategy progressively enhances out-of-distribution performance.
Inference. Given that the model primarily predicts relative quality rankings between video pairs, a practical conversion to absolute quality scores is required. To achieve this, we adopt an adaptive soft comparison method proposed in  [77], which first computes a soft probability matrix across ranking categories by comparing the test video against anchor videos, followed by maximum a posteriori (MAP) [54] estimation under Thurstone’s Case V model [53] to obtain calibrated quality scores.
For anchor selection, the range of quality scores of PLVD-VQA is partitioned into five quality intervals. Within each interval, the video exhibiting the lowest inter-model score variance is selected as the anchor, ensuring stable reference points for score derivation. We list the detailed calculation procedure in the supplement file.
5 Experiments
5.1 Experimental Setups
Implementation Details For each stage of training, we initialize the image encoder, image projector, and LLM with LLaVA-ov-chat (7B) weights [24] and the motion encoder with SlowFast pre-trained weights. The model is trained for one epoch on four NVIDIA A800 GPUs with a total batch size of $4$ and a learning rate of 11e-55. Inference is performed on two NVIDIA RTX 3090 GPUs.
Benchmark Datasets The proposed methods is evaluated on ten VQA benchmarks, categorized into in-domain and out-of-distribution (OOD) groups to systematically assess their generalization. The in-domain evaluation benchmarks include LSVQ Test [70], LSVQ 1080p [70], KoNViD-1k [16], LIVE-VQC [50], and YouTube-UGC [60], all of which contain UGC videos. The OOD datasets consist of LIVE-YT-Gaming [71], CGVDS [47], LIVE-YT-HFR [33], Waterloo-IVC-4K [26], and KVQ [32]. Specifically, LIVE-YT-Gaming and CGVDS contain gaming videos, which exhibit significant differences in content compared to natural UGC videos. LIVE-YT-HFR includes videos with varying frame rates, focusing on temporal-related distortions. Waterloo-IVC-4K and KVQ primarily evaluate compression artifacts, where Waterloo-IVC-4K contains compressed high-resolution videos at different spatial resolutions, while KVQ comprises compressed UGC videos processed with various video enhancement techniques. We introduce details of these datasets in the supplement file.
Evaluation Criteria We adopt two widely used criteria to evaluate the performance of VQA models: Spearman Rank Correlation (SRCC) and Pearson Linear Correlation (PLCC), which indicate the prediction monotonicity and prediction linearity, respectively.
Competing Methods We compare our model with six opinion-unaware BVQA methods, including NIQE [38], IL-NIQE [72], VIIDEO [39], STEM [20], TPQI [27], and BUONA-VISTA [63], as well as five supervised VQA methods: Minimalistic-VQA (VII) [52], Minimalistic-VQA (IX) [52], FAST-VQA [61], DOVER [64], and Q-Align [65]. These five supervised VQA models serve as strong baselines, as they are used to pseudo-label video quality and can be regarded as teacher models. The supervised models are trained on the LSVQ training set.
Table 3: Supervised performance comparison of our model against competitive methods. The experimental results on KoNViD-1k, YouTube-UGC, and LIVE-VQC are obtained using 10-fold cross-validation. “Base” denotes training LMM-PVQA without self-supervised pretraining
Testing Set $_{\text{test}} LSVQ$$_{\text{1080p}}$ KoNViD-1k LIVE-VQC YouTube-UGC
TLVQM [22] 0.772 / 0.774 0.589 / 0.616 0.732 / 0.724 0.670 / 0.691 0.685 / 0.692
VIDEAL [55] 0.794 / 0.783 0.545 / 0.554 0.751 / 0.741 0.630 /0.640 0.687 / 0.709
VSFA [25] 0.801 / 0.796 0.675 / 0.704 0.794 / 0.799 0.718 / 0.771 0.787 / 0.789
PatchVQ [70] 0.827 / 0.828 0.711 / 0.739 0.791 / 0.786 0.827 / 0.837 NA / NA
SimpleVQA [51] 0.867 / 0.861 0.764 / 0.803 0.856 / 0.860 0.845 / 0.859 0.847 / 0.856
FAST-VQA [61] 0.880 / 0.880 0.781 / 0.813 0.891 / 0.892 0.849 / 0.865 0.855 / 0.852
DOVER [64] 0.878 / 0.866 0.782 / 0.813 0.908 / 0.910 0.860 / 0.875 0.841 / 0.851
MinimalisticVQA [52] 0.881 / 0.879 0.781 / 0.820 0.889 / 0.890 0.842 / 0.854 0.890 / 0.891
Soft Ranking (Base) 0.880 / 0.864 0.790 / 0.814 0.896 / 0.877 0.840 / 0.853 0.858 / 0.848
Soft Ranking (Stage 1) 0.907 / 0.904 0.832 / 0.857 0.911 / 0.908 0.884 / 0.894 0.910 / 0.911
Table 4: Experimental results of ablation study. “Spatial” and “Motion” denote the spatial branch and the temporal branch of the visual feature extractor. In Stage 1, if PLVD-DS (Part1) is not specified, the model is trained solely on PLVD-VQA (Part1). In Stage 2, “w/o label refinement” and “w/ label refinement” indicate whether the trained model is used to label the data in PLVD-Part2
Stage 1 Stage 2 $_{\text{test}} LSVQ$$_{\text{1080p}}$ KoNViD-1k LIVE-VQC YouTube-UGC Overall
Spatial Motion PLVD-DS w/o label refinement w/ label refinement SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
0.872 0.862 0.787 0.814 0.868 0.864 0.783 0.819 0.834 0.839 0.843 0.846
0.883 0.872 0.804 0.827 0.883 0.876 0.799 0.830 0.842 0.843 0.855 0.857
0.886 0.880 0.803 0.830 0.891 0.888 0.797 0.832 0.845 0.849 0.858 0.863
0.887 0.879 0.801 0.829 0.890 0.884 0.791 0.826 0.853 0.855 0.858 0.862
0.887 0.880 0.802 0.830 0.893 0.892 0.798 0.833 0.856 0.859 0.859 0.864
Stage 1 Stage 2 LIVE-YT-Gaming CGVDS LIVE-YT-HFR Waterloo-IVC-4K KVQ Overall
Spatial Motion PLVD-DS w/o label refinement w/ label refinement SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
0.678 0.735 0.754 0.800 0.334 0.413 0.404 0.459 0.616 0.656 0.561 0.610
0.688 0.736 0.769 0.808 0.446 0.477 0.425 0.470 0.619 0.662 0.579 0.622
0.697 0.752 0.799 0.829 0.481 0.525 0.552 0.614 0.690 0.725 0.650 0.693
0.688 0.741 0.794 0.822 0.446 0.508 0.541 0.608 0.705 0.735 0.651 0.694
0.717 0.763 0.810 0.834 0.515 0.611 0.633 0.696 0.749 0.764 0.704 0.741
5.2 Performance Analysis
The performance of all compared methods and our proposed models is summarized in Table 2, with detailed analysis conducted from the following perspectives:
Hard Ranking v.s. Soft Ranking. The model trained with soft ranking marginally outperforms its hard-ranking counterpart on in-domain VQA benchmarks and CGVDS. Moreover, it achieves significantly better performance on Waterloo-IVC-4K and KVQ, while exhibiting slightly inferior results on LIVE-YT-Gaming and LIVE-YT-HFR. Overall, the soft ranking strategy incorporates “quality distance” information into the training process, enabling the model to easily learn quality comparison capabilities and achieve better performance.
Iterative Self-Improvement Training Strategy. Consistent performance improvements are observed across both in-domain and OOD datasets as the model advances through training stages. This empirical evidence demonstrates that our iterative training strategy effectively enhances model capability through progressive self-learning. Notably, substantial gains are achieved on challenging VQA benchmarks where existing models struggle: the LIVE-YT-HFR, Waterloo-IVC-4K, and KVQ datasets exhibit relative improvements of $16.2\%$, $21.4\%$, $9.3\%$ in SRCC respectively, after three-stage refinement.
In-Domain Performance Comparison. We observe that all proposed model variants outperform the five competing VQA models, which are used for label annotation in Section 3.2.1, confirming our self-supervised approach successfully distills ensemble knowledge from its teacher models on unlabeled data. However, our models are slightly outperformed by FAST-VQA and DOVER on LIVE-VQC, which exhibits more complex temporal distortions compared to other datasets [52]. We attribute this performance gap to architectural differences: FAST-VQA and DOVER fine-tune a 3D DNN (i.e., VideoSwin) specifically optimized for spatiotemporal feature extraction, whereas our framework handles temporal information using a pre-trained SlowFast model.
OOD Performance Comparison. Our models significantly outperform competing approaches on the OOD benchmarks, achieving a $24.7\%$ improvement in SRCC for overall OOD evaluation compared to the competing approaches. We attribute the strong OOD performance to two key factors. First, our training datasets include compression distortions (H.264 and H.265), and the synthetic quality ranking labels provide useful supervision for assessing compression-related video quality for these datasets. Second, our iterative self-improvement training strategy enables progressive adaptation to unseen distortion types, such as frame rate inconsistencies in LIVE-YT-HFR, AVS/VP9 compression artifacts in Waterloo-IVC-4K, and enhancement-induced distortions in KVQ—none of which are present in the training data. These results demonstrate the effectiveness of the proposed self-supervised VQA framework in real-world quality assessment scenarios.
5.3 Supervised Performance Comparison
We conduct supervised fine-tuning of our self-supervised trained model on the in-domain VQA benchmarks, as shown in Table 3. The results reveal that: (1) Our fine-tuned model surpasses all competing VQA baselines with 3.5% higher SRCC, and (2) it achieves $6.4\%$ improvement over the baseline trained without self-supervised initialization. These results confirm that self-supervised representation learning critically enhances downstream fine-tuning effectiveness.
Furthermore, we compare the prediction accuracy of pairwise comparisons in LMM-PVQA with four open-source LMMs that have demonstrated strong performance in video understanding. As shown in Table 6, LMM-PVQA significantly outperforms these advanced LMMs, providing highly accurate quantitative predictions. In contrast, other models primarily rely on high-level instruction-tuning datasets, resulting in suboptimal accuracy in quality prediction. This suggests that existing LMMs exhibit limited quality perception capabilities, highlighting the advantages of our proposed approach in video quality assessment.
5.4 Ablation Study
We conduct the ablation experiments to validate the effectiveness of LMM-PVQA, with results shown in Table 4.
Motion Encoder. Incorporating the motion encoder for temporal distortion representation yields consistent improvements across all benchmarks, particularly achieving $33.5\%$ SRCC improvement on LIVE-YT-HFR—the benchmark specifically designed for high frame-rate distortion analysis. This verifies that explicit motion modeling is essential for capturing temporal-related distortions.
Synthetic Distortion Data. The incorporation of the PLVD-DS dataset also demonstrates consistent performance improvements across all benchmarks, yielding marginal gains on in-domain benchmarks while achieving substantial enhancements on OOD benchmarks. It can be attributed that the PLVD-VQA dataset have already equipped the model with quality assessment capabilities for in-domain content, whereas the PLVD-DS dataset help mitigate the domain gap on OOD benchmarks by simulating common artificial degradations (e.g., compression).
Iterative Self-Improvement Training Strategy. Since our iterative self-improvement training strategy introduces new training samples, it is critical to verify that performance gains stemming from the strategy itself rather than additional data. To isolate this effect, we train a Stage 2 model using the same base training set (PLVD-Part1/Part2), where videos in PLVD-Part2 are labeled by five VQA judges, without incorporating the Stage 1 model as the new judge. Experimental results confirm that merely increasing training data (without iterative refinement) fails to surpass the Stage 1 model’s performance, thereby validating the effectiveness of our strategy in enabling self-improvement through iterative optimization.
6 Conclusion
We propose a self-supervised learning framework for VQA to mitigate dependence on human-annotated datasets. Our approach constructs a large-scale video pair dataset labeled via established VQA models and synthetic distortion simulations, then adopts a learning-to-rank paradigm to learn quality assessment capabilities. By integrating an iterative self-improvement training strategy, our model progressively enhances its evaluation performance through self-learning. Our method achieves state-of-the-art zero-shot performance on both in-domain and OOD benchmarks, demonstrating its effectiveness.
Limitations. Our method is inherently designed for pairwise quality comparison. To obtain absolute quality scores for a single video, it must be compared against multiple anchor videos (five in this study), which significantly increases inference time.
Future Work. We plan to explore more automated video pair annotation strategies, such as leveraging expert-domain VQA models (e.g., VMAF for video compression), utilizing LMMs with carefully designed prompt engineering, and employing text-to-video generation algorithms to synthesize videos of varying quality through specified prompts, thus substantially expanding the diversity and scale of training data. Furthermore, we intend to extend our framework to incorporate additional modalities, such as images and audio, to develop a more generalizable quality assessment model.
Supplementary Material
Figure 4: Examples videos of different categories in our large dataset.
Appendix A More Details of Our PLVD Database
A.1 Analysis of the Collected Videos
As shown in Fig. 5, our dataset is collected from multiple popular social media platforms with relatively uniform sampling, comprising $20\%$ from Bilibili, $20\%$ from Youku, $25\%$ from YouTube, and $35\%$ from TikTok. Notably, our dataset covers a diverse range of content categories, exceeding twenty in total. In addition to common categories such as lifestyle, food, and animals, it also includes specialized categories such as gaming, AI-generated content, and high-resolution content. To illustrate the diversity of our dataset, we present a variety of video samples in Fig. 4, showcasing the broad range of content available in our large-scale video quality assessment (VQA) dataset. Unlike existing datasets, which often focus on specific formats, our dataset encompasses a wider variety of formats, including both landscape and portrait orientations, as well as various resolutions. This diversity enhances the comprehensiveness of our dataset, making it more suitable for evaluating video quality across a wide kinds of scenarios.
Figure 5: Our dataset is collected from multiple popular social media platforms and encompasses a wide range of content categories.
A.2 Analysis of Low-level Metrics
Our data selection strategy is based on a mixed-integer programming method [56], which optimizes dataset composition by aligning feature histograms. Specifically, we utilize this approach to match the distributions of nine low-level metrics (blockiness [45], blur [41], contrast [43], noise, flickering [42], colourfulness [14], luminance, spatial information (SI) [19], and temporal information (TI) [19]) between our dataset and widely used VQA datasets. Each metric is computed as follows:
Blockiness
[45] is quantified by analyzing the luminance differences between pixels within and across encoding blocks. Specifically, we compute the absolute luminance differences between adjacent pixel pairs within the same encoding block (internal pixel pairs) and those spanning adjacent blocks (external pixel pairs). The blockiness metric is then determined as the ratio of the total sum of internal pixel difference values to the total sum of external pixel difference values across the entire video frame:
$B=\frac{\sum_{(x,y)\in\mathcal{I}}|I(x,y)-I(x+1,y)|}{\sum_{(x,y)\in\mathcal{E}}|I(x,y)-I(x+1,y)|},$ (3)
where $I(x,y)$ represents the luminance value at pixel location $(x,y)$, $\mathcal{I}$ denotes the set of internal pixel pairs, and $\mathcal{E}$ represents the set of external pixel pairs. A higher blockiness value indicates stronger blocking artifacts, which typically result from aggressive video compression.
Blur
is measured using the Cumulative Probability of Blur Detection (CPBD) [41], which evaluates perceptual sharpness based on edge width distribution. A higher CPBD value indicates a sharper image. Given an edge pixel $e_{i}$, its width $w(e_{i})$ is compared with the Just Noticeale Blur (JNB) threshold, determining the blur detection probability $w_{JNB}(e_{i})$. The final CPBD score is computed as:
$\text{CPBD}=P(P_{\text{BLUR}}\leq P_{\text{JNB}})=\sum_{P_{\text{BLUR}}=0}^{P_{\text{JNB}}}P(P_{\text{BLUR}}).$ (4)
Contrast
is a measure of the dispersion of pixel intensity values within the video frame and can be quantified using the standard deviation of grayscale intensities [43]. Specifically, for a grayscale image $I(x,y)$, the mean intensity $\mu$ is first computed as:
$\mu=\frac{1}{M\times N}\sum_{x=1}^{M}\sum_{y=1}^{N}I(x,y),$ (5)
where $M$ and $N$ denote the width and height of the image, respectively, and $I(x,y)$ represents the intensity at pixel $(x,y)$. The contrast value $\sigma$ is then obtained by calculating the standard deviation of intensity values:
$\sigma=\sqrt{\frac{1}{M\times N}\sum_{x=1}^{M}\sum_{y=1}^{N}(I(x,y)-\mu)^{2}}.$ (6)
The standard deviation $\sigma$ represents the contrast of the video frame, where a higher $\sigma$ value indicates a greater dispersion of intensity values and thus a higher contrast.
Flickering
occurs when an encoder skips macroblocks to conserve bitrate, especially in low-texture, slow-motion regions [42]. It is quantified by counting macroblock transitions from an “unupdated” to an “updated” state, with a threshold $T_{f}$ ensuring only significant changes are considered. The flickering metric is computed as:
$F=\frac{1}{M\times N}\sum_{x=1}^{M}\sum_{y=1}^{N}\mathbb{I}\left(|I_{t}(x,y)-I_{t-1}(x,y)|>T_{f}\right),$ (7)
where $I_{t}(x,y)$ is the luminance at pixel $(x,y)$ in frame $t$, and $\mathbb{I}(\cdot)$ is an indicator function. A higher $F$ indicates stronger flickering artifacts.
Colourfulness
quantifies color distribution differences across RGB channels, following [14]. Given a frame with RGB channels $R,G,B$, we compute:
$r_{g}=R-G,\quad y_{b}=\frac{1}{2}(R+G)-B.$ (8)
The Colourfulness metric is then:
$C=\sqrt{\sigma_{r_{g}}^{2}+\sigma_{y_{b}}^{2}}+0.3\times\sqrt{\mu_{r_{g}}^{2}+\mu_{y_{b}}^{2}},$ (9)
where $\sigma$ and $\mu$ denote the standard deviations and means of $r_{g}$ and $y_{b}$, respectively.
Luminance
is measured as the combined intensity of the three RGB channels, defined as:
$L=R+G+B.$ (10)
SI
measures spatial complexity using the Sobel filter. The standard deviation of the Sobel-filtered frame over all pixels is computed, and the maximum value over time represents the SI:
$SI=\max_{time}\left\{\text{std}_{space}\left[\text{Sobel}(F_{n})\right]\right\}.$ (11)
Figure 6: Illustration of different levels of spatial distortion video frames in our large-scale dataset.
TI
measures motion intensity by calculating the difference between consecutive frames. The temporal difference at pixel $(i,j)$ is:
$M_{n}(i,j)=F_{n}(i,j)-F_{n-1}(i,j).$ (12)
The TI value is the maximum standard deviation of $M_{n}(i,j)$ over time and space:
$TI=\max_{time}\left\{\text{std}_{space}[M_{n}(i,j)]\right\}.$ (13)
To optimize computational efficiency, all metrics are extracted at a sampling rate of one frame per second.
A.3 More Details of PLVD-DS Datas
Figure 7: Illustration of different levels of streaming distortion video frames in our large-scale dataset.
A.3.1 Spatial Distortion
We introduce five common spatial distortions: resizing, Gaussian blur, Gaussian noise, darkening, and brightening. Each distortion is applied at five different levels to simulate varying degrees of degradation, ranging from mild to severe. Fig. 6 illustrates examples of these distortions, where the quality of video frames progressively deteriorates as the distortion level increases. Below, we provide details on how these spatial distortions are generated, where $I$ represents the original frame, and $I^{\prime}$ denotes the distorted frame.
Resizing:
The frame is first downsampled by a scaling factor $s$ and then upsampled back to its original size. This process reduces spatial details and introduces pixelation artifacts, simulating resolution loss. The transformation is defined as:
$I^{\prime}=\text{Upsample}(\text{Downsample}(I,s),s),$ (14)
where $s$ takes values from the set $\{2,3,4,8,16\}$.
Gaussian Blur:
The frame is convolved with a Gaussian kernel, where the standard deviation $\sigma_{blur}$ controls the extent of the blur. A larger $\sigma_{blur}$ results in a wider spread of the Gaussian function, leading to a stronger blurring effect by averaging pixel intensities over a larger neighborhood. The blurring process is defined as:
$I^{\prime}=I*G(\sigma_{blur}),$ (15)
where $G(\sigma_{blur})$ is a Gaussian kernel with standard deviation $\sigma_{blur}$ which takes values from the set $\{0.1,0.5,1,2,5\}$, and * denotes the convolution operation.
Gaussian noise:
Gaussian noise is introduced by adding random variations to each pixel, following a normal distribution with mean $\mu$ and standard deviation $\sigma_{noise}$. The noise level is controlled by adjusting $\sigma_{noise}$, where higher values result in more pronounced noise artifacts. The process is defined as:
$I^{\prime}=I+N(\mu,\sigma_{noise}^{2}),$ (16)
where $N(\mu,\sigma_{noise}^{2})$ represents Gaussian noise with mean $\mu$ and variance $\sigma_{noise}^{2}$, added independently to each pixel. $\sigma$ takes values from the set $\{0.001,0.002,0.003,0.005,0.01\}$.
Darkening:
Darkening is applied by reducing the luminance component in the color space. The effect is controlled by a parameter $p$, which determines the degree of brightness reduction. The luminance channel $L$ is adjusted using an interpolation function $f(L,p)$ as follows:
$L^{\prime}=f(L,p).$ (17)
The parameter $p$ is selected from a predefined set of values $\{0.05,0.1,0.2,0.4,0.8\}$, with larger values leading to stronger darkening effects.
Table 5: An overview of our testing datasets
Dataset Year # of Videos # of Scenes Resolution Duration Frame Rate Distortion Type
KoNViD-1k [16] 2017 1,200 1,200 540p 8 24, 25, 30 In-the-wild
LIVE-VQC [50] 2018 585 585 240p–1080p 10 30 In-the-wild
YouTube-UGC [60] 2019 1,380 1,380 360p–4K 20 30 In-the-wild
LSVQ [70] 2021 38,811 38,811 99p–4K 5–12 << 60 In-the-wild
Waterloo-IVC-4K [26] 2019 1200 20 540p, 1080p, 4k 9-10 24, 25, 30 H.264 compression
LIVE-YT-HFR [33] 2021 480 16 1080p 6-10 24, 30, 60, 82, 98, 120 Frame rate, VP9 compression
LIVE-YT-Gaming [71] 2022 600 600 360p–1080p 8–9 30, 60 PGC, UGC
CGVDS [47] 2023 360 15 480p, 720p, 1080p 30 20, 30, 60 H.264 compression
KVQ [32] 2024 4200 600 - 3-8 - UGC
Brightening:
In contrast, brightening is achieved by enhancing the luminance component in the color space. The luminance channel $L$ is modified using a nonlinear transformation function $g(L,p)$:
$L^{\prime}=g(L,p),$ (18)
The parameter $p$ is selected from $\{0.1,0.2,0.4,0.7,1.1\}$, with larger values producing a stronger brightening effects.
A.3.2 Temporal Distortion
We introduce two types of temporal distortions: jitter and stuttering, each distortion maintain three different levels.
Jitter:
Jitter introduces random shifts and random cropping followed by resizing of video frames. The amount of shift is determined by the jitter level, which controls the extent of spatial displacement.
For each frame, random horizontal and vertical shifts are applied using an affine transformation matrix, which shifts the frame along the $x$- and $y$-axes. Additionally, each frame is cropped by a small amount from the edges and resized back to its original dimensions, simulating pixelation effects or lower-quality views. The transformation matrix is described as follows:
$M=\begin{bmatrix}1&0&\text{random\_shift\_x}\\ 0&1&\text{random\_shift\_y}\end{bmatrix}$ (19)
where random_shift_x and random_shift_y are random values determined by the jitter level.
Stuttering:
Stuttering is introduced by randomly dropping frames at a controlled rate. The drop rate $p_{d}$ is determined by the distortion level, where higher levels correspond to increased frame loss. For each frame $I_{t}$, a random probability is drawn and compared with $p_{d}$. If the frame is dropped, it is replaced by the previous frame $I_{t-1}$, simulating temporal freezing in the video. The process can be formulated as:
$I_{t}^{\prime}=\begin{cases}I_{t-1},&\text{if }r<p_{d},\\ I_{t},&\text{otherwise}\end{cases}$ (20)
where $r\sim U(0,1)$ is a random variable drawn from a uniform distribution.
A.3.3 Streaming Distortion
As illustrated in Fig. 7, we select the two most common compression standards, H.264 and H.265, to simulate video quality degradation for the compression distortion. These distortions are applied using the ffmpeg tool, a widely used multimedia framework, to encode the videos with different compression settings. Specifically, we chose four fixed constant rate factor (CRF) values for each compression standard to control the level of distortion.
For H.264 compression, we selected the fast encoding mode, which provides a good balance between encoding speed and compression efficiency, making it suitable for real-time applications. To cover a wide range of compression levels, we applied H.264 compression using CRF values of 24, 36, 48, and 63, ensuring the simulation of various quality degradation scenarios.
In contrast, for H.265 compression, we selected the very slow encoding mode, which prioritizes compression efficiency over speed, leading to higher quality video at the cost of longer encoding times. To achieve fine-grained quality simulation, we applied H.265 compression with a narrower CRF range of 36, 40, 44, and 48, allowing for precise control over compression artifacts.
These encoding settings help to simulate typical real-world compression scenarios, where different modes and CRF values are chosen based on the trade-off between video quality and encoding performance.
A.4 More Details on Testing Datasets
Table 5 provides an overview of our testing datasets, which encompass diverse content types, resolutions, durations, frame rates, and distortion types. The first four datasets consist of in-the-wild videos containing various authentic distortions, while the remaining datasets focus on specific content types and distortion factors. For example, LIVE-YT-Gaming is dedicated to gaming content, LIVE-YT-HFR targets frame rate distortions, and Waterloo-IVC-4K covers different types of compression artifacts. By evaluating our model across these nine datasets, we demonstrate its robustness and effectiveness in both in-domain and out-of-distribution (OOD) quality assessment scenarios.
Table 6: Performance comparison of our model in terms of prediction accuracy on four VQA datasets against advanced LLMs. Additionally, performance comparison of an ensemble of five methods versus a single model on the same datasets
Method $_{\textbf{test}} LSVQ$$_{\textbf{1080p}}$ KoNViD-1k LIVE-VQC YouTube-UGC
LLaVA-NeXT-Video (8B) [74] 0.508 0.531 0.512 0.531 0.506
mPLUG-Owl3-8B [68] 0.547 0.534 0.605 0.548 0.543
InternVL2_\_5-8B [59] 0.566 0.629 0.583 0.596 0.588
Qwen2.5-7B-VL [2] 0.696 0.666 0.740 0.664 0.631
LLaVA-ov-chat [24] 0.667 0.641 0.727 0.664 0.639
MinimalisticVQA(VII) [52] 0.907 0.817 0.906 0.839 0.784
MinimalisticVQA(IX)  [52] 0.926 0.845 0.904 0.843 0.820
FAST-VQA [61] 0.918 0.836 0.918 0.832 0.765
DOVER [64] 0.922 0.840 0.915 0.860 0.812
Q-Align [65] 0.922 0.849 0.923 0.846 0.822
Ensemble five methods 0.939 0.864 0.935 0.860 0.831
Soft Ranking (Stage 1) 0.926 0.855 0.935 0.849 0.853
Appendix B More Details of Pairwise Quality Annotation
B.1 VQA Models for Pseudo-labeling
We choose five SOTA VQA models: Minimalistic-VQA (VII) [52], Minimalistic-VQA (IX) [52], FAST-VQA [61], DOVER [64], and Q-Align [65] as our initial judges to formulate our pairwise quality annotation. As shown in Table 6, the ranking accuracy of the five ensemble methods is higher than that of any single model. The detail introduction of the five models is as follows:
Minimalistic-VQA (VII)
employs Swin Transformer-B [31], pre-trained on ImageNet-1K [9], as the spatial quality analyzer to extract quality-aware spatial features from key frames, ensuring robust spatial quality assessment.
Minimalistic-VQA (IX)
builds upon Minimalistic-VQA (VII) by incorporating a temporal quality analyzer to account for motion distortions. The temporal quality analyzer, implemented using the SlowFast [12] network pre-trained on the Kinetics-400 [5] dataset, extracts motion-related features from video chunks, enhancing the model’s ability to assess temporal quality variations.
FAST-VQA
introduces Grid Mini-patch Sampling (GMS) strategy, which preserves local quality by sampling patches at raw resolution and maintains global quality through uniformly sampled mini-patches. These mini-patches are spliced and temporally aligned into fragments. To process these fragments, the Fragment Attention Network (FANet) is designed to effectively extract video quality features. Combining GMS and FANet, FAST-VQA achieves efficient end-to-end video quality assessment with effective feature representation learning.
DOVER
builds upon FAST-VQA as its technical branch to capture low-level distortions, while introducing an additional aesthetic branch to assess high-level semantic composition, which relates to user preferences and content recommendation. By disentangling these two perspectives, DOVER establishes a more human-aligned and interpretable framework for video quality assessment.
Q-Align
presents a novel training strategy for large multimodal model (LMM) in VQA by replacing direct numerical score predictions with discrete, text-defined rating levels (e.g., “excellent”, “good”, “fair”, “poor”, “bad”) as learning targets. During inference, Q-Align extracts the log probabilities of each rating level, applies softmax normalization to obtain a probability distribution, and computes a weighted average to derive the final predicted quality score.
B.2 Prompts for Model Training
We final construct the label prompt of our large dataset using a fixed template:
Question: "Now you will receive two videos. The first video:<image>. The second video:<image>. Please watch these videos carefully, and then answer the following question: Comparing with the first video, how do you assess the quality of the second video?"
Answer: "The quality of the second video is [level] to the first video."
Appendix C More Details of Our Model
C.1 Training Details
The model is trained using the DeepSpeed framework with mixed-precision floating-point operations to optimize memory and computational efficiency. The training is conducted for one epoch with a batch size of 1 per device and a gradient accumulation step of 1. The optimizer follows AdamW with a learning rate of $1\times 10^{-5}$, a cosine learning rate schedule, and a warmup ratio of 0.03.
We employ a joint training strategy for images and videos. For the image encoder, videos are sampled at a rate of one frame per second, with each sampled frame resized to a resolution of $384\times 384$, while images are directly resized to the same resolution. For the motion encoder, videos are fully encoded across all frames to capture temporal dynamics, whereas images, which lack temporal information, are assigned an all-zero tensor as their temporal representation.
C.2 Inferring Details
C.2.1 Probability Modeling
Though we employ video pairs to train our model by enabling it to determine whether the second video is better than the first, our goal during inference is to obtain an absolute quality score for a single video. To achieve this, we propose a method that converts the probability of a test video being better or worse than anchor videos into a final quality score.
First, we describe how to construct the probability distribution for comparative quality assessments. For hard ranking, the comparative token set is defined as:
$\mathcal{S}=\{s_{k}\}_{k=1}^{2}=\{\textit{worse},\textit{better}\}.$ (21)
For soft ranking, the comparative token set is extended to:
$\mathcal{S}=\{s_{k}\}_{k=1}^{5}=\{\textit{inferior},\textit{worse},\textit{similar},\textit{better},\textit{superior}\}.$ (22)
The probability of each token is computed using the softmax function:
$q_{s_{k}}=\frac{e^{s_{k}}}{\sum_{m=1}^{r}e^{s_{m}}},$ (23)
where $q_{s_{k}}$ represents the probability of the $k$-th token, and $r$ denotes the number of levels.
To obtain a quality score for the test video $v_{\text{eval}}$, we aggregate its comparative probabilities against anchor videos using a weighted summation:
$P\left(v_{\text{anchor}},v_{\text{eval}}\right)=\sum_{k=1}^{r}\alpha_{k}q_{s_{k}}\left(v_{\text{anchor}},v_{\text{eval}}\right),\quad r=1\dots p.$ (24)
where $\alpha_{k}$ are fixed weights that reflect the comparative levels. Specifically, for hard ranking, the weights are:
$\{\alpha_{k}\}_{k=1}^{2}=\{0,0.5\}.$ (25)
For soft ranking, the weights are defined as:
$\{\alpha_{k}\}_{k=1}^{5}=\{0,0.25,0.5,0.75,1\}.$ (26)
This approach enables the model to generate a continuous quality score for a single video by leveraging its relative comparisons against anchor videos in the training set.
C.2.2 Score Modeling
Finally, we construct a probability matrix based on pairwise comparisons with a set of anchor videos. Given a set of five anchor videos, we first define a probability matrix:
$M_{r}\in\mathbb{R}^{5\times 5},$ (27)
where each entry $P(b^{(i)},b^{(j)})$ represents the probability that anchor video $b^{(i)}$ is preferred over $b^{(j)}$. This probability satisfies:
$P(b^{(i)},b^{(j)})=1-P(b^{(j)},b^{(i)}),\quad P(b^{(i)},b^{(i)})=0.5.$ (28)
To evaluate a test video $v_{\text{test}}$, we compute its comparative probabilities against all anchor videos, forming the probability vector:
$c=\left[P(b^{(1)},v_{\text{test}}),P(b^{(2)},v_{\text{test}}),\dots,P(b^{(5)},v_{\text{test}})\right].$ (29)
Next, we integrate this vector into the complete probability matrix:
$M\in\mathbb{R}^{(5+1)\times(5+1)},M=\begin{bmatrix}M_{r}&c\\ (1-c)^{\top}&0.5\end{bmatrix}$. (30)
With this probability matrix, we estimate the final quality score using maximum a posteriori (MAP) [54] estimation under Thurstone’s Case V model [53]. This is formulated as the following convex optimization problem:
$\arg\max_{\hat{q}}$ $\sum_{i,j}M_{i,j}\log\left(\Phi(\hat{q}^{(i)}-\hat{q}^{(j)})\right)$ (31)
$-\sum_{i}\frac{\hat{q}^{(i)}}{2},\quad\text{s.t.}\sum_{i}\hat{q}^{(i)}=0.$
Here, $\Phi(\cdot)$ denotes the standard normal cumulative distribution function, and the final score $\hat{q}^{(n+1)}$ corresponds to the estimated quality of the test video.
Appendix D More Details of Experimental Results
We also compare the prediction accuracy of our model with several advanced LLMs, including LLaVA-NeXT-Video (7B) [74], mPLUG-Owl3-8B [68], InternVL2.5-8B [59], Qwen2.5-7B-VL [2], and our base model LLaVA-ov-chat [24], as shown in Table 6. Evidently, our model significantly outperforms all LLMs, suggesting that existing LLMs trained on high-level tasks still struggle with low-level visual perception.